A cannon is mounted on a railroad car. The cannon fires a 1.75kg ball with a muzzle velocity 300.m/s to the left. The mass of the cannon-car combo is 4500.kg. if all is initially at rest, what is the recoil velocity of the cannon-car?

1
Expert's answer
2022-02-19T04:56:10-0500

Given:

"m_{cannon}=4500\:\rm kg"

"m_{ball}=1.75\:\rm kg"

"v_{ball}=300\:\rm m/s"


The law of conservation of momentum says


"m_{ball}v_{ball}=m_{cannon}v_{cannon}"

Hence, the recoil speed of the gun


"v_{cannon}=\frac{m_{ball}v_{ball}}{m_{cannon}}=\frac{1.75*300}{4500}=0.12\:\rm m/s"